OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variables. (x+4)(x+2) / 4x+8 Answers: 1. X+2/4 (x= -2) 2. x+4/4 (x = -2) 3. 4/x+4 (x =-4) 4. 4/x+2 (x = -2)

OpenStudy (hockeychick23):

ok so to simplify this you need to factor the bottom of the given equation: (x+4)(x+2) / 4x+8 --> (x+4)(x+2)/4(x+2) Now you can cancel out x+2 on the top and bottom. You are now left with x+4/4

OpenStudy (hockeychick23):

(your answer is 2)

OpenStudy (anonymous):

thank you! Would you mind helping with a couple more?

OpenStudy (hockeychick23):

sure

OpenStudy (anonymous):

Simplify and select the answer with the appropriate restrictions for the variables. 2x-3x^2/5x^3 That's the next one, i'll get you the answers

OpenStudy (anonymous):

1. 1/5 2. 1/5x (x = 0) 3. 1/5x^2 (x = 0) 4. 2-3x/5x^2 (x = 0)

OpenStudy (hockeychick23):

ok so the top of this equation is factored out to be: -x(3x-2) the bottom of this equation cannot be factored. x does= 0

OpenStudy (anonymous):

So that means answer one would be taken out of the possibilities. But there's three left and nothing else can be factored out of it?

OpenStudy (anonymous):

Nevermind, i got it.